HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

INTERACTIONS OF ATMOSPHERIC POLLUTANTS WITH WATER: 4-METHYL-2-NITROPHENOL AND ITS WATER COMPLEX STUDIED BY

MICROWAVE SPECTROSCOPY AND QUANTUM CHEMISTRY

\underline{E. ANTONELLI} ,^{1} S. BAWEJA,^{2} S. HUSSAIN,^{2} A. FERNANDEZ-RAMOS,^{3} I. KLEINER,^{4} H. V. L. NGUYEN,^{1,5} M. E. SANZ^{2},^{1}Univ Paris Est Créteil and Université Paris Cité, CNRS, LISA, F-94010 Créteil, France,^{2}Department of

Chemistry, King’s College London, Britannia House, 7 Trinity Street, London SE1 1DB, UK,^{3}Departamento de Química Física and Centro Singular de Investigación en Química Biolóxica e Materiais Moleculares (CIQUS), Jenaro de la Fuente s/n, Universidad de Santiago de Compostela, 15782 Santiago de Compostela, Spain, 4Université Paris Cité and Univ Paris Est Créteil, CNRS, LISA, F-75013 Paris,

France,^{5}Institut Universitaire de France (IUF), F-75231 Paris, France

The microwave spectrum of 4-methyl-2-nitrophenol (4MNP) and its complex with one water molecule was measured by a molecular jet chirped-pulse Fourier transform microwave (FTMW) spectrometer working in the frequency range from 2 to 8 GHz and for the monomer, a resonator FTMW spectrometer from 2 to 26.5 GHz in addition. Quantum chemical calculations were performed at the B3LYP-D3BJ/6-311++G(d,p) and MP2/6-311++G(d,p) levels of theory to obtain optimized molecular geometries. One conformer for 4MNP and two different isomers for 4MNP-H_{2}O were identified in the rotational spectrum. Large tunneling splittings arising from the low barrier to internal rotation of the methyl group and hyperfine structures from the quadrupole coupling of the ^{14}N nucleus were observed for all species. The spectra were analyzed and fitted using the XIAM^{1} and BELGI-C_{s}-hyperfine^{2} codes. The standard deviations of the fits achieve measurement accuracy of 4 kHz for the monomer and 10 kHz for the water complexes. The deduced V_{3} potential values of 106 cm^{−1} for the monomer and the more stable water complex as well as 158 cm^{−1} for the second water complex are in reasonable agreement with the values predicted by quantum chemistry.

1doi:10.1515/zna-1996-0807, H. Hartwig, H. Dreizler, Z. Naturforsch., 51, 923-932 (1996).

2doi:10.1021/acs.jpca.6b02111, R. Kannengießer, W. Stahl, H.V.L. Nguyen, I. Kleiner, J. Phys.

Chem. A, 120, 3992-3997 (2016).

p-number: p158  Submitted on Wed, 14 Jun 2023 22:17:40 +0200